# LexEval: A Comprehensive Chinese Legal Benchmark for Evaluating Large Language Models

### v1

@@ 4.2 Evaluated Models
GPT-4 27.2 34.8 14.0 79.8 51.0 94.0 77.2 96.2 79.2 68.3 53.7 33.2 66.0 57.8
Qwen-14B-Chat 28.0 38.6 11.4 93.4 45.3 90.0 85.6 91.8 79.5 90.2 67.4 31.6 44.7 48.6
Qwen-7B-Chat 22.8 38.9 8.4 79.8 43.3 87.0 67.2 92.0 79.2 83.9 25.8 24.2 36.3 29.6
ChatGPT 19.0 25.6 7.0 56.8 42.3 87.0 76.0 82.2 77.7 60.3 23.0 19.4 39.6 38.2
@@
GPT-4 34.0 39.1 26.9 14.2 38.9 15.7 65.2 55.2 75.8 52.1 1
Qwen-14B-Chat 30.0 31.9 33.4 24.1 35.7 18.6 29.2 42.0 63.0 50.2 2
Qwen-7B-Chat 21.0 28.6 31.5 19.2 34.7 18.3 22.1 38.9 56.8 43.0 3
ChatGPT 28.4 22.4 24.0 10.7 34.8 16.8 33.7 32.1 55.8 39.7 4
@@
ChatGLM2 19.8 21.1 28.3 15.0 24.1 14.0 36.8 27.2 52.2 29.8 9
Chinese-Alpaca-2-7B 27.8 24.7 28.6 15.7 31.2 14.6 21.5 28.4 40.4 29.4 10
Fuzi-Mingcha 20.0 16.1 54.0 20.7 21.4 17.4 10.8 13.1 25.0 29.2 11
@@
XVERSE-13B 10.4 12.2 12.1 13.9 6.8 19.0 19.9 29.4 55.0 28.0 14
Ziya-LLaMA-13B 2.2 20.7 24.2 15.3 33.3 17.9 9.6 18.6 20.4 27.3 15
@@ 4.3 Experimental Results
Qwen-14B-Chat 34.0 46.8 17.1 95.6 32.7 89.0 89.6 85.6 77.3 83.4 70.6 31.2 43.4 43.2
Qwen-7B-Chat 23.0 41.5 8.7 82.4 32.7 83.0 60.4 50.0 78.4 79.0 50.4 24.0 36.5 28.6
@@ 5 Conclusion & Future Work
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix B.1.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Appendix B.2.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Section 3.4.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Appendix A.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] We don’t have a training process.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Because running a massive amount of LLMs is expensive.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix E.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 3.2.
    2. (b)
      Did you mention the license of the assets? [Yes] See Appendix C.1.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] See Appendix A.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix C.1.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Appendix 3.4.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.2, Appendix C.4 and Appendix F.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Section 3.2.
@@ Appendix F Guidelines for Expert-Annotation
MoSS-Moon-sft 21.0 18.8 22.3 20.1 28.2 15.1 7.8 18.4 22.6 22.8 24
Baichuan-7B-base 16.0 14.5 21.8 30.4 16.6 9.8 11.0 18.7 35.4 22.8 25
LLaMA-2-7B-Chat 7.4 18.4 18.8 7.6 11.2 14.0 8.4 11.9 19.2 22.2 26
InternLM-7B 9.2 9.9 20.5 28.8 5.1 9.5 16.7 23.9 42.6 21.4 27
LLaMA-2-7B 18.6 13.2 29.4 9.4 6.0 11.5 5.0 14.7 16.2 18.4 28
@@
[…]
MoSS-Moon-sft 25.0 18.1 7.9 13.9 25.2 9.4 11.1 12.1 24.2 $5\uparrow 2$0
[…]
Baichuan-7B-base 3.2 19.4 1.5 11.5 1.4 9.6 8.4 19.9 22.2 $3\uparrow 2$6

### v2

@@ 4.2 Evaluated Models
GPT-4 34.0 34.8 14.0 79.8 51.0 94.0 77.2 96.2 79.2 68.3 53.7 33.2 66.0 57.8
Qwen-14B-Chat 28.0 38.6 11.4 93.4 45.3 90.0 85.6 91.8 79.5 90.2 27.9 31.6 44.7 50.4
Qwen-7B-Chat 22.8 38.9 8.4 79.8 43.3 87.0 67.2 92.0 79.2 83.9 53.2 24.2 36.3 45.0
ChatGPT 19.0 25.6 9.0 56.8 42.3 87.0 76.0 82.2 77.7 60.3 23.0 19.4 39.6 38.2
@@
GPT-4 33.8 39.1 25.0 16.0 38.3 13.6 65.2 55.2 75.8 52.2 1
Qwen-14B-Chat 30.0 31.9 33.9 23.1 36.0 19.1 29.2 42.0 63.0 48.5 2
Qwen-7B-Chat 21.0 28.6 30.8 19.0 34.7 18.3 22.1 38.9 56.8 44.8 3
ChatGPT 28.4 22.0 22.8 13.1 34.3 13.1 33.7 32.1 55.8 39.6 4
@@
Fuzi-Mingcha 20.0 16.1 57.8 27.8 21.4 17.3 10.8 13.1 25.0 30.2 9
ChatGLM2 19.8 21.1 28.3 15.0 24.1 14.0 36.8 27.2 52.2 29.8 10
Chinese-Alpaca-2-7B 27.8 24.7 28.6 15.7 31.2 14.6 21.5 28.4 40.4 29.4 11
@@
XVERSE-13B 10.4 12.2 12.1 13.9 6.8 19.0 19.9 29.4 55.0 27.7 14
TigerBot-base 25.8 23.0 20.8 11.3 34.5 12.6 16.3 19.0 39.2 27.3 15
@@ 4.3 Experimental Results
Qwen-14B-Chat 34.0 49.7 13.4 95.4 42.7 92.0 88.4 88.2 58.6 90.7 61.2 34.2 47.2 41.0
Qwen-7B-Chat 23.2 42.3 8.7 82.2 34.7 85.0 60.4 49.2 78.1 77.2 61.6 24.4 35.8 41.8
@@ Appendix F Guidelines for Expert-Annotation
Baichuan-7B-base 16.0 14.5 21.8 30.4 16.6 9.8 11.0 18.7 35.4 22.8 24
LLaMA-2-7B-Chat 7.4 18.4 18.8 7.6 11.2 14.0 8.4 11.9 19.2 22.2 25
InternLM-7B 9.2 9.9 20.5 28.8 5.1 9.5 16.7 23.9 42.6 21.4 26
LLaMA-2-7B 18.6 13.2 29.4 9.4 6.0 11.5 5.0 14.7 16.2 18.4 27
ChatLaw-13B 26.8 10.5 23.5 13.4 27.4 14.0 13.4 8.1 14.2 17.9 28
@@
Lawyer-LLaMA 0.4 12.5 16.4 13.3 22.9 15.7 20.3 25.1 32.6 $5\uparrow 2$0